OpenStudy (anonymous):

Solve the equation:

OpenStudy (anonymous):

\[\frac{ 9x }{ 7 }-x=\frac{ x-28 }{ 21 }\]

OpenStudy (anonymous):

add x to both sides..

OpenStudy (anonymous):

9x/7 = (x-28)/7 + x ....nw solve make deno same..))

OpenStudy (lgbasallote):

interesting...

OpenStudy (lgbasallote):

btw..im referring to @Yahoo! 's reply not your question @GenGen96 ...i totally have no idea about your question..we havent discussed it in class yet

OpenStudy (anonymous):

Hahaha, okay.

OpenStudy (anonymous):

Don't SPAM

OpenStudy (anonymous):

i will complain..))

OpenStudy (lgbasallote):

heh. gengen is still thinking about your hint. you might as well kill time

OpenStudy (anonymous):

I"m confused, though. I'm still unsure as to how to solve it from there.

OpenStudy (lgbasallote):

you're in good hands with @Yahoo! ...he's one of the best here. and i mean that with no sarcasm :D

OpenStudy (anonymous):

hm...hm.....u r teasing me..))

OpenStudy (unklerhaukus):

\[\frac{ 9x }{ 7 }-x=\frac{ x-28 }{ 21 }\] \[\frac{ 9x }{ 7 }=\frac{ x-28 }{ 21 }+x\] now multiply everything by twenty one

OpenStudy (lgbasallote):

yes @Yahoo! i think that so much of me and that so little of you

OpenStudy (anonymous):

Everything? Like the 9x/7 and x-28/21+x combined?

OpenStudy (anonymous):

@lgbasallote i dont knw to write in LATeX...so..it will be messy

OpenStudy (unklerhaukus):

\[21\times\frac{ 9x }{ 7 }=21\times\frac{ x-28 }{ 21 }+21\times x\]

OpenStudy (lgbasallote):

you're supposed to be yahoo. no wonder everyone uses google now

OpenStudy (anonymous):

Thanks!

OpenStudy (lgbasallote):

i guess my words helped her...?

OpenStudy (anonymous):

u r gonna get banned soon...))) before u reach SS100

OpenStudy (anonymous):

banned forever*

OpenStudy (lgbasallote):

i'd like that. i'd reaaally like that.